Problem 1-11
The Tax Formula for Individuals, Filing Status and Tax Computation, The Standard Deduction (LO 1.3, 1.5, 1.7)

Christine is a single 50-year-old taxpayer with no dependents. Her only income is $40,750 of wages. Calculate her taxable income and her tax liability.

Table for the standard deduction

Filing Status Standard Deduction
Single $ 12,200
Married, filing jointly 24,400
Married, filing separately 12,200
Head of household 18,350
Qualifying widow(er) 24,400

Click here to access the tax tables.

Taxable income: ?

Tax liability: ?

$
Answer #1
2019 federal income tax brackets
Calculation of total income and taxliability
Total Income 40750
Filing Status Amount
Single
Total Income 40750
Less:Standard Deduction -12200
Taxable income 28550
Tax Amount              = (28550-9700)*12%+(9700)*10%
3232
Married, filing jointly
Total Income 40750
Less:Standard Deduction -24400
Taxable income 16350
Tax Amount              = 16350*10%
1635
Married, filing separately
Total Income 40750
Less:Standard Deduction -12200
Taxable income 28550
Tax Amount              = (28550-9700)*12%+(9700)*10%
3232
Head of household
Total Income 40750
Less:Standard Deduction -18350
Taxable income 22400
Tax Amount              = (22400-13850)*12%+(13850)*10%
2411
Qualifying Widow (or Qualifying Widower)
Total Income 40750
Less:Standard Deduction -24400
Taxable income 16350
Tax Amount              = 16350*10%
1635
Qualifying Widow (or Qualifying Widower) is a filing status that allows you to retain the benefits of the Married Filing Jointly status for two years after the year of your spouse's death.